suppose the length of a side of a cube is x^10 find the volume of a cube in terms of x
Answer by Alan3354(69443)

suppose the length of a side of a cube is x^10 find the volume of a cube in terms of x
--------
Vol = s^3 = (x^10)^3
Vol = x^30